# Principal stratification with $U$-statistics under principal ignorability

### v2

@@ 2.3 Identification assumptions
Assumption 1 (Treatment ignorability).
@@
Assumption 2 states that the treatment assignment only affects the intermediate outcome in a non-negative (or non-positive if assume $D_{i}(1)\leq D_{i}(0)$) direction, and restricts the number of possible strata by excluding the harmed or defiers stratum $S=01$ (or $S=10$). For example, when $D$ represents the survival status, monotonicity holds when the treatment is expected not to harm the survival of each unit. When $D$ represents actual treatment received in noncompliance problems, monotonicity holds if units in the control condition have no access to active treatment.
Assumption 3 (Mean principal ignorability).
@@
Assumption 3 rules out unmeasured confounding between the principal strata membership and the final outcome, ensuring that the PGCE components can be directly estimated from the observed data. For example, under Assumptions 1-3, the conditional expectations in group $\mathcal{S}_{10}$ are all equal to $\mathbb{E}[h\{Y_{1}(1),Y_{2}(0)\}|Z_{1}=1,D_{1}=1,Z_{2}=0,D_{2}=0,X_{1},X_{2}]$. In other words, within the observed data stratum defined by $(Z_{1}=1,D_{1}=1,Z_{2}=0,D_{2}=0)$, conditional on pre-treatment covariates, the combinations of possible latent principal strata, $(S_{1}=10,S_{2}=10)$, $(S_{1}=11,S_{2}=10)$, $(S_{1}=10,S_{2}=00)$, and $(S_{1}=11,S_{2}=00)$, are exchangeable with respect to the expected contrast. Similarly, the conditional expected contrasts within group $\mathcal{S}_{11}$ and $\mathcal{S}_{00}$ are all equal to estimable conditional expectations, $\mathbb{E}[h\{Y_{1}(1),Y_{2}(0)\}|Z_{1}=1,D_{1}=1,Z_{2}=0,D_{2}=2,X_{1},X_{2}]$ and $\mathbb{E}[h\{Y_{1}(1),Y_{2}(0)\}|Z_{1}=1,D_{1}=0,Z_{2}=0,D_{2}=0,X_{1},X_{2}]$, respectively. That is, with Assumption 3, each one of the subpopulations, $(Z_{1}=1,D_{1}=1,Z_{2}=0,D_{2}=0)$, $(Z_{1}=1,D_{1}=1,Z_{2}=0,D_{2}=1)$, and $(Z_{1}=1,D_{1}=0,Z_{2}=0,D_{2}=0)$, can be viewed as mixtures of its corresponding subgroups defined by observed covariates only.
Remark 2.1.
@@
Assumptions 2 and 3 are instrumental in achieving nonparametric identification of the PGCE, however, their applicability, in general, cannot be directly tested using observed data alone without prior knowledge. Although our primary focus is on identifying and efficiently estimating the PGCE under these two structural assumptions, we provide a confounding function approach for sensitivity analyses under assumed violations of these two structural assumptions in Section S7 of the Online Supplement.
@@ 3.1 Identification of the PGCE estimands
To obtain nonparametric identification formulas for the PGCE estimands, we start by defining components of the observed data distribution, i.e., the treatment probability, the principal score, and the mean of the outcome contrast conditional on covariates (referred to as pairwise outcome mean hereafter for brevity). We consider both the observational and experimental studies, and define the treatment probability or the propensity score, by $\pi(X)=\mathbb{P}(Z=1|X)$, which is the probability of a participant receiving the treatment conditional on pre-treatment covariates (Rosenbaum and Rubin,, 1983). In addition, the principal score, $e_{s}(X)=\mathbb{P}(S=s|X)$, is defined as the conditional probability that a participant belongs to the principal stratum $s$ given the covariates (Ding and Lu,, 2016), for $s\in\{10,00,11\}$ under Assumption 2. Furthermore, we use averages over the distribution of covariates, $\pi=\mathbb{E}\{\pi(X)\}=\mathbb{P}(Z=1)$ and $e_{s}=\mathbb{E}\{e_{s}(X)\}=\mathbb{P}(S=s)$, to denote population fractions of receiving treatment $z$ or belonging to stratum $s$, respectively. Lastly, we define the pairwise outcome mean within the observed stratum $(Z_{1}=z_{1},D_{1}=d_{1},Z_{2}=z_{2},D_{2}=d_{2})$, as
@@
for $z_{1},z_{2},d_{1},d_{2}\in\{0,1\}$. We provide some examples of the pairwise outcome mean models for specific contrast functions below.
Example 3.1 (Pairwise mean function for average treatment effect).
@@
Example 3.2 (Pairwise mean function for probability index).
For studying the probability index estimand, the contrast function is specified as $h(u,v)=\mathbbm{1}(u\geq v)$. If the potential outcomes are continuous, a Gaussian regression model can be considered to represent the mean potential outcomes as in Example 3.1. In this case, following Wolfe and Hogg, (1971), one can derive the pairwise outcome mean as
@@
Example 3.3 (Pairwise outcome mean for win estimands).
Consider the win estimands developed in Example 2.2, where the contrast function is given by the vector $h(u,v)=\{\mathbbm{1}(u>v),\mathbbm{1}(u<v)\}$ (the win estimands are then up to transformations of the two-dimensional estimand). For ordinal outcomes, if we specify the proportional odds or ordinal logistic regression model for $Y(z)$ conditional on $D$ and $X$, Mao, (2017) has shown that the resulting pairwise outcome mean functions are given by the two-dimensional vector
@@
For estimating the principal scores, we first define conditional probability, $p_{z}(X)=\mathbb{P}(D=1|Z=z,X)$, with population fraction $p_{z}=\mathbb{E}\{p_{z}(X)\}=\mathbb{P}(D=1|Z=z)$. Then under Assumption 2, the principal scores can be expressed as (Jiang et al.,, 2022)
@@
This is because, with monotonicity, the observed stratum $(Z=0,D=1)$ only contains stratum $S=11$, and observed stratum $(Z=1,D=0)$ only contains stratum $S=00$. The remaining principal score for stratum $S=10$ can thus be obtained by subtracting $e_{11}(X)$ and $e_{00}(X)$ from one. The population fractions of the three principal strata can be similarly connected as $e_{10}=p_{1}-p_{0}$, $e_{00}=1-p_{1}$, and $e_{11}=p_{0}$. We will refer to $\{p_{0}(X),p_{1}(X)\}$ as principal scores due to the one-to-one mapping to $\{e_{10}(X),e_{00}(X),e_{11}(X)\}$. These elements then prepare us to state the set of nonparametric identification results in the following Theorem 3.4.
Theorem 3.4 (Nonparametric identification).
@@
We use $\tau_{h}^{10}$ as an example to elaborate on the intuitions behind Theorem 3.4, and the identification of the other two PGCEs follows the same strategy. In Theorem 3.4(a), $\tau_{h}^{10}$ is expressed as a doubly inverse probability-weighted average of the observed contrast function; the essential part of the weighting scheme leverages the principal scores to tilt all pairs of $(\mathcal{O}_{i},\mathcal{O}_{j})$ ($i\neq j$) in the observed stratum $(Z_{1}=1,D_{1}=1,Z_{2}=0,D_{2}=0)$ to represent the covariate distribution in the principal stratum $S=10$. Specifically, under Assumptions 1 and 2, we have
@@
which is proportional to the conditional probability of each pair belonging to the principal stratum $S=10$, and by the Bayes’ Theorem, is equivalent to the density ratio of $(X_{1},X_{2})$ conditional and unconditional on $S_{1}=S_{2}=10$. This ensures the identification of $\tau_{h}^{10}$. In Theorem 3.4(b), while $\tau_{h}^{10}$ is expressed using the treatment probability and the pairwise outcome mean, it can be shown that the conditional expectation of the weight attached to the pairwise outcome mean, under Assumptions 1 and 2, is,
@@
Therefore, $\tau_{h}^{10}$ is identified as the pairwise outcome mean function averaged over the density $f(X_{1},X_{2}|S_{1}=S_{2}=10)/f(X_{1},X_{2})$. Finally, in Theorem 3.4(c), $\tau_{h}^{10}$ is expressed using the principal score and pairwise outcome mean, where, in comparison with Theorem 3.4(b), the treatment probability and intermediate outcome, $DZ/\pi(X)-D(1-Z)/\{1-\pi(X)\}$, is replaced by the principal score, $p_{1}(X)-p_{0}(X)$. The conditional expectation of the weight on covariates attached to the pairwise outcome mean, under Assumptions 1 and 2, remains $e_{10}(X_{1})e_{10}(X_{2})/e_{10}^{2}$, and hence $\tau_{h}^{10}$ is identified as an average of the pairwise outcome mean function over the target covariate density.
Remark 3.1.
Theorem 3.4 is a natural extension of the results for nonparametric identification of the principal average causal effects (Jiang et al.,, 2022, Theorem 1) to the PCGEs based on an arbitrary contrast function $h$; and it reduces to the Theorem 1 in Jiang et al., (2022) when $h(u,v)=u-v$ with $\mu_{z_{1}d_{1}z_{2}d_{2}}(X_{1},X_{2})=\mu_{z_{1}d_{1}}(X_{1})-\mu_{z_{2}d_{2}}(X_{2})$.
@@ 3.2 Estimators based on nonparametric identification formulas
For point estimation, a common practice for operationalizing the nonparametric identification formulas is to assume parametric models for nuisance functions, i.e., $\pi(X;\alpha)$, $p_{z}(X;\beta)$ or $e_{s}(X;\beta)$, and $\mu_{z_{1}d_{1}z_{2}d_{2}}(X_{1},X_{2};\gamma)$, where $\alpha$, $\beta$, and $\gamma$ are finite-dimensional nuisance parameters, with their corresponding true values denoted by $\alpha^{*}$, $\beta^{*}$, and $\gamma^{*}$. We write the maximum likelihood estimators for these finite-dimensional parameters as $\widehat{\alpha}$, $\widehat{\beta}$, and $\widehat{\gamma}$. To aid further discussions of statistical properties of estimators for the PGCEs, we introduce notations $\mathcal{M}_{\pi}$, $\mathcal{M}_{p}$, and $\mathcal{M}_{\mu}$ to denote conditions where the treatment probability, the principal score, and the outcome mean are correctly specified, respectively. Specifically, under $\mathcal{M}_{\pi}$, we have $\pi(X;\alpha^{*})=\pi(X)$; under $\mathcal{M}_{p}$, we have $p_{z}(X;\beta^{*})=p_{z}(X)$ and $e_{s}(X;\beta^{*})=e_{s}(X)$ for $z\in\{0,1\}$ and $s\in\{10,00,11\}$; and under $\mathcal{M}_{\mu}$, we have $\mu_{z_{1}d_{1}z_{2}d_{2}}(X_{1},X_{2};\gamma^{*})=\mu_{z_{1}d_{1}z_{2}d_{2}}(X_{1},X_{2})$ for $z_{1},z_{2},d_{1},d_{2}\in\{0,1\}$. Additionally, we use ‘+’ in the subscript to denote conditions where more than one nuisance function is correctly specified. In addition, the union notation, $\cup$’, denotes the condition where at least one nuisance function is correctly specified.
We can obtain plug-in estimators by replacing components in the identification formulas with their empirical counterparts, and expectations with empirical averages. For convenience, we first introduce the following notations: $\mathbb{E}_{n}f(\mathcal{O})=n^{-1}\sum_{i=1}^{n}f(\mathcal{O}_{i})$ for any $f(\mathcal{O})$, and $\mathbb{U}_{n}f(\mathcal{O}_{1},\mathcal{O}_{2})=\binom{n}{2}^{-1}\sum_{i<j}f(\mathcal{O}_{i},\mathcal{O}_{j})$ for any $f(\mathcal{O}_{1},\mathcal{O}_{2})$. Along the lines in Jiang et al., (2022), we define the following quantity for any function $f(Y,D,X)$:
@@
Equation (11) is useful in the estimation of $p_{z}$, as it leads to a doubly robust estimator once we treat the intermediate variable as the outcome (Tsiatis,, 2006):
@@
Plug-in estimators for the PGCEs take the form of $U$-statistics (van der Vaart,, 1998). For example, Theorem 3.4(a) motivates the following estimator for $\tau_{h}^{10}$ based on the estimated treatment probability and the principal score:
@@
Theorem 3.4(b) motivates the following estimator for $\tau_{h}^{10}$ based on the estimated propensity score and the pairwise outcome mean:
@@
and Theorem 3.4(c) motivates the following estimator for $\tau_{h}^{10}$ based on the estimated principal score and the pairwise outcome mean:
@@
Remark 3.2.
The weighting estimator $\widehat{\tau}_{h,\pi+p}^{s}$ leverages the inverse propensity score and the principal score, which is an extension of the inverse propensity score weighted estimator in Mao, (2017) for estimating $\tau_{h}$ in the absence of the intermediate variable. The theoretical basis for the consistency and asymptotic normality of $\widehat{\tau}_{h,\pi+p}^{s}$, $\widehat{\tau}_{h,\pi+\mu}^{s}$, and $\widehat{\tau}_{h,p+\mu}^{s}$, can be established by first recognizing their ratio structures and then investigating the numerators and the denominators separately after applying the Taylor’s expansion on the ratio estimators (Jiang et al.,, 2022); see the Online Supplement for additional details. The asymptotic behavior for the denominators involving $\widehat{p}_{z}$ can be straightforwardly obtained using empirical process theory for the parametric class (van der Vaart,, 1998, §19.2). The numerators are parameter-indexed $U$-processes whose asymptotic properties can be established following Lemma S1 and Proposition 1 of Mao, (2017).

### v3

@@ 2.3 Identification assumptions
Assumption 1 (Treatment ignorability)
@@
Assumption 2 states that the treatment assignment only affects the intermediate outcome in a non-negative (or non-positive if assume $D_{i}(1)\leq D_{i}(0)$) direction and restricts the number of possible strata by excluding the harmed or defiers stratum $S=01$ (or $S=10$). For example, when $D$ represents the survival status, monotonicity holds when the treatment is expected not to harm the survival of each unit. When $D$ represents actual treatment received in noncompliance problems, monotonicity holds if units in the control condition have no access to active treatment.
Assumption 3 (Mean principal ignorability)
@@
Assumption 3 rules out unmeasured confounding between the principal strata membership and the final outcome, ensuring that the PGCE components can be directly estimated from the observed data. For example, under Assumptions 1-3, the conditional expectations in group $\mathcal{S}_{10}$ are all equal to $\mathbb{E}[h\{Y_{1}(1),Y_{2}(0)\}|Z_{1}=1,D_{1}=1,Z_{2}=0,D_{2}=0,X_{1},X_{2}]$. In other words, within the observed data stratum defined by $(Z_{1}=1,D_{1}=1,Z_{2}=0,D_{2}=0)$, conditional on pre-treatment covariates, the combinations of possible latent principal strata, $(S_{1}=10,S_{2}=10)$, $(S_{1}=11,S_{2}=10)$, $(S_{1}=10,S_{2}=00)$, and $(S_{1}=11,S_{2}=00)$, are exchangeable with respect to the expected contrast. Similarly, the conditional expected contrasts within group $\mathcal{S}_{11}$ and $\mathcal{S}_{00}$ are all equal to estimable conditional expectations, $\mathbb{E}[h\{Y_{1}(1),Y_{2}(0)\}|Z_{1}=1,D_{1}=1,Z_{2}=0,D_{2}=1,X_{1},X_{2}]$ and $\mathbb{E}[h\{Y_{1}(1),Y_{2}(0)\}|Z_{1}=1,D_{1}=0,Z_{2}=0,D_{2}=0,X_{1},X_{2}]$, respectively. That is, with Assumption 3, each one of the subpopulations, $(Z_{1}=1,D_{1}=1,Z_{2}=0,D_{2}=0)$, $(Z_{1}=1,D_{1}=1,Z_{2}=0,D_{2}=1)$, and $(Z_{1}=1,D_{1}=0,Z_{2}=0,D_{2}=0)$, can be viewed as mixtures of its corresponding subgroups defined by observed covariates only.
Remark 1
@@
Assumptions 2 and 3 are instrumental in achieving nonparametric identification of the PGCE; however, their applicability, in general, cannot be directly tested using observed data alone without prior knowledge. Although our primary focus is identifying and efficiently estimating the PGCE under these two structural assumptions, we provide a confounding function approach for sensitivity analyses under assumed violations of these two structural assumptions in Section S7 of the Online Supplement.
@@ 3.1 Identification of the PGCE estimands
To obtain nonparametric identification formulas for the PGCE estimands, we start by defining components of the observed data distribution, i.e., the treatment probability, the principal score, and the mean of the outcome contrast conditional on covariates (referred to as pairwise outcome mean hereafter for brevity). We consider both the observational and experimental studies and define the treatment probability or the propensity score, by $\pi(X)=\mathbb{P}(Z=1|X)$, which is the probability of a participant receiving the treatment conditional on pre-treatment covariates (Rosenbaum and Rubin, 1983). In addition, the principal score, $e_{s}(X)=\mathbb{P}(S=s|X)$, is defined as the conditional probability that a participant belongs to the principal stratum $s$ given the covariates (Ding and Lu, 2016), for $s\in\{10,00,11\}$ under Assumption 2. Furthermore, we use averages over the distribution of covariates, $\pi=\mathbb{E}\{\pi(X)\}=\mathbb{P}(Z=1)$ and $e_{s}=\mathbb{E}\{e_{s}(X)\}=\mathbb{P}(S=s)$, to denote population fractions of receiving treatment $z$ or belonging to stratum $s$, respectively. Lastly, we define the pairwise outcome mean within the observed stratum $(Z_{1}=z_{1},D_{1}=d_{1},Z_{2}=z_{2},D_{2}=d_{2})$, as
@@
for $z_{1},z_{2},d_{1},d_{2}\in\{0,1\}$. Below, we provide some examples of the pairwise outcome mean models for specific contrast functions.
Example 3 (Pairwise mean function for average treatment effect)
@@
Example 4 (Pairwise mean function for probability index)
For studying the probability index estimand, the contrast function is specified as $h(u,v)=\mathbbm{1}(u\geq v)$. If the potential outcomes are continuous, a Gaussian regression model can be considered to represent the mean potential outcomes as in Example 3. In this case, following Wolfe and Hogg (1971), one can derive the pairwise outcome mean as
@@
Example 5 (Pairwise outcome mean for win estimands)
Consider the win estimands developed in Example 2, where the contrast function is given by the vector $h(u,v)=\{\mathbbm{1}(u>v),\mathbbm{1}(u<v)\}$ (the win estimands are then up to transformations of the two-dimensional estimand). For ordinal outcomes, if we specify the proportional odds or ordinal logistic regression model for $Y(z)$ conditional on $D$ and $X$, Mao (2017) has shown that the resulting pairwise outcome mean functions are given by the two-dimensional vector
@@
For estimating the principal scores, we first define the conditional probability, $p_{z}(X)=\mathbb{P}(D=1|Z=z,X)$, with population fraction $p_{z}=\mathbb{E}\{p_{z}(X)\}=\mathbb{P}(D=1|Z=z)$. Then under Assumption 2, the principal scores can be expressed as (Jiang et al., 2022)
@@
This is because, with monotonicity, the observed stratum $(Z=0,D=1)$ only contains stratum $S=11$, and observed stratum $(Z=1,D=0)$ only contains stratum $S=00$. The remaining principal score for stratum $S=10$ can thus be obtained by subtracting $e_{11}(X)$ and $e_{00}(X)$ from one. The population fractions of the three principal strata can be similarly connected as $e_{10}=p_{1}-p_{0}$, $e_{00}=1-p_{1}$, and $e_{11}=p_{0}$. We will refer to $\{p_{0}(X),p_{1}(X)\}$ as principal scores due to the one-to-one mapping to $\{e_{10}(X),e_{00}(X),e_{11}(X)\}$. These elements then prepare us to state the set of nonparametric identification results in the following Theorem 1.
Theorem 1 (Nonparametric identification)
@@
We use $\tau_{h}^{10}$ as an example to elaborate on the intuitions behind Theorem 1, and the identification of the other two PGCEs follows the same strategy. In Theorem 1(a), $\tau_{h}^{10}$ is expressed as a doubly inverse probability-weighted average of the observed contrast function; the essential part of the weighting scheme leverages the principal scores to tilt all pairs of $(\mathcal{O}_{i},\mathcal{O}_{j})$ ($i\neq j$) in the observed stratum $(Z_{1}=1,D_{1}=1,Z_{2}=0,D_{2}=0)$ to represent the covariate distribution in the principal stratum $S=10$. Specifically, under Assumptions 1 and 2, we have
@@
which is proportional to the conditional probability of each pair belonging to the principal stratum $S=10$, and by Bayes’ Theorem, is equivalent to the density ratio of $(X_{1},X_{2})$ conditional and unconditional on $S_{1}=S_{2}=10$. This ensures the identification of $\tau_{h}^{10}$. In Theorem 1(b), while $\tau_{h}^{10}$ is expressed using the treatment probability and the pairwise outcome mean, it can be shown that the conditional expectation of the weight attached to the pairwise outcome mean, under Assumptions 1 and 2, is,
@@
Therefore, $\tau_{h}^{10}$ is identified as the pairwise outcome mean function averaged over the density $f(X_{1},X_{2}|S_{1}=S_{2}=10)/f(X_{1},X_{2})$. Finally, in Theorem 1(c), $\tau_{h}^{10}$ is expressed using the principal score and pairwise outcome mean, where, in comparison with Theorem 1(b), the treatment probability and intermediate outcome, $DZ/\pi(X)-D(1-Z)/\{1-\pi(X)\}$, is replaced by the principal score, $p_{1}(X)-p_{0}(X)$. The conditional expectation of the weight on covariates attached to the pairwise outcome mean, under Assumptions 1 and 2, remains $e_{10}(X_{1})e_{10}(X_{2})/e_{10}^{2}$, and hence $\tau_{h}^{10}$ is identified as an average of the pairwise outcome mean function over the target covariate density.
Remark 2
Theorem 1 is a natural extension of the results for nonparametric identification of the principal average causal effects (Jiang et al., 2022, Theorem 1) to the PCGEs based on an arbitrary contrast function $h$; and it reduces to the Theorem 1 in Jiang et al. (2022) when $h(u,v)=u-v$ with $\mu_{z_{1}d_{1}z_{2}d_{2}}(X_{1},X_{2})=\mu_{z_{1}d_{1}}(X_{1})-\mu_{z_{2}d_{2}}(X_{2})$.
@@ 3.2 Estimators based on nonparametric identification formulas
For point estimation, a common practice for operationalizing the nonparametric identification formulas is to assume parametric models for nuisance functions, i.e., $\pi(X;\alpha)$, $p_{z}(X;\beta)$ or $e_{s}(X;\beta)$, and $\mu_{z_{1}d_{1}z_{2}d_{2}}(X_{1},X_{2};\gamma)$, where $\alpha$, $\beta$, and $\gamma$ are finite-dimensional nuisance parameters, with their corresponding true values denoted by $\alpha^{*}$, $\beta^{*}$, and $\gamma^{*}$. We write the maximum likelihood estimators for these finite-dimensional parameters as $\widehat{\alpha}$, $\widehat{\beta}$, and $\widehat{\gamma}$. To aid further discussions of statistical properties of estimators for the PGCEs, we introduce notations $\mathcal{M}_{\pi}$, $\mathcal{M}_{p}$, and $\mathcal{M}_{\mu}$ to denote conditions where the treatment probability, the principal score, and the outcome mean are correctly specified, respectively. Specifically, under $\mathcal{M}_{\pi}$, we have $\pi(X;\alpha^{*})=\pi(X)$; under $\mathcal{M}_{p}$, we have $p_{z}(X;\beta^{*})=p_{z}(X)$ and $e_{s}(X;\beta^{*})=e_{s}(X)$ for $z\in\{0,1\}$ and $s\in\{10,00,11\}$; and under $\mathcal{M}_{\mu}$, we have $\mu_{z_{1}d_{1}z_{2}d_{2}}(X_{1},X_{2};\gamma^{*})=\mu_{z_{1}d_{1}z_{2}d_{2}}(X_{1},X_{2})$ for $z_{1},z_{2},d_{1},d_{2}\in\{0,1\}$. Additionally, we use ‘+’ in the subscript to denote conditions where more than one nuisance function is correctly specified, and $\cup$’ to denote the condition where at least one nuisance function is correctly specified.
We can obtain plug-in estimators by replacing components in the identification formulas with their empirical counterparts and expectations with empirical averages. For convenience, we first introduce the following notations: $\mathbb{E}_{n}f(\mathcal{O})=n^{-1}\sum_{i=1}^{n}f(\mathcal{O}_{i})$ for any $f(\mathcal{O})$, and $\mathbb{U}_{n}f(\mathcal{O}_{1},\mathcal{O}_{2})=\binom{n}{2}^{-1}\sum_{i<j}f(\mathcal{O}_{i},\mathcal{O}_{j})$ for any $f(\mathcal{O}_{1},\mathcal{O}_{2})$. Along the lines in Jiang et al. (2022), we define the following quantity for any function $f(Y,D,X)$:
@@
Equation (11) is useful in the estimation of $p_{z}$, as it leads to a doubly robust estimator once we treat the intermediate variable as the outcome (Tsiatis, 2006):
@@
Plug-in estimators for the PGCEs take the form of $U$-statistics (van der Vaart, 1998). For example, Theorem 1(a) motivates the following estimator for $\tau_{h}^{10}$ based on the estimated treatment probability and the principal score:
@@
Theorem 1(b) motivates the following estimator for $\tau_{h}^{10}$ based on the estimated propensity score and the pairwise outcome mean:
@@
and Theorem 1(c) motivates the following estimator for $\tau_{h}^{10}$ based on the estimated principal score and the pairwise outcome mean:
@@
Remark 3
The weighting estimator $\widehat{\tau}_{h,\pi+p}^{s}$ leverages the inverse propensity score and the principal score, which is an extension of the inverse propensity score weighted estimator in Mao (2017) for estimating $\tau_{h}$ in the absence of the intermediate variable. The theoretical basis for the consistency and asymptotic normality of $\widehat{\tau}_{h,\pi+p}^{s}$, $\widehat{\tau}_{h,\pi+\mu}^{s}$, and $\widehat{\tau}_{h,p+\mu}^{s}$, can be established by first recognizing their ratio structures and then investigating the numerators and the denominators separately after applying the Taylor’s expansion on the ratio estimators (Jiang et al., 2022); see the Online Supplement for additional details. The asymptotic behavior for the denominators involving $\widehat{p}_{z}$ can be straightforwardly obtained using empirical process theory for the parametric class (van der Vaart, 1998, §19.2). The numerators are parameter-indexed $U$-processes whose asymptotic properties can be established following Lemma S1 and Proposition 1 of Mao (2017).